OpenStudy (anonymous):

Write the equation of the line that passes through (2, 2) and (6, 3) in standard form? I ended up with fractions... anyone mind walking me through it ?

OpenStudy (cwrw238):

slope = 3-2 / 6-2) = 1/4 use the standard form y - y1 = m(x - x1) m=slope and (x1,y1) is a point on the line so use the point (2,2): y - 2 = (1/4)(x - 2) now multiply the equation by 4 to get rid of fraction: 4y - 8 = x - 2 x - 4y = -6 - standard form

OpenStudy (anonymous):

thank you! But how did you get your ending result answer from 4y - 8 = x - 2?

OpenStudy (cwrw238):

ok? -need further help?

OpenStudy (cwrw238):

ok - i did some steps in my head 4y - 8 = x - 2 add 8 to both sides 4y = x + 6 subtract 4y form both sides 0 = x - 4y + 6 now subtract 6 from both sides -6 = x - 4y then flip the values over the = to give x - 4y = -6 as i understand it standard form is like ax + by = c

OpenStudy (anonymous):

now i see;) thank you!